OpenStudy (anonymous):

Find the equation of the line that is tangent to the function y=-4sinx at x= pi/4

OpenStudy (anonymous):

\[\huge x=\frac{ \pi }{ 4 } \]

OpenStudy (anonymous):

\[\huge y'= -2\sqrt{2}\]

OpenStudy (anonymous):

There are two steps you need to complete here. First find the derivative, and then plug in the value of x into that derivative. Do you know how to take the derivative?

OpenStudy (anonymous):

Wait why is the derivative -2(sqrt2)?

OpenStudy (anonymous):

\[\huge y'(\frac{ \pi }{ 4 })\]

OpenStudy (anonymous):

I forgot to but the brackets

OpenStudy (anonymous):

Oh okay. Well just do as I said originally and find the derivative. Do you know what it is?

OpenStudy (anonymous):

\[\huge y'= -4\cos(\frac{ \pi }{ 4 })\]

OpenStudy (anonymous):

So the slope of the tangent line is going to the value you get when you plug in x. You already seemed to have gotten it, so that's the first part. I'm actually blanking out on finding the y intercept of that line... I apologize, it's been a while -.-

OpenStudy (anonymous):

It's okay lol :P

OpenStudy (anonymous):

find the derivative of the y function and then plug in the x point to find the slope of the tangent line

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i need the equation though @yummydum

OpenStudy (e.mccormick):

You have the slope and a point. How about the point-slope formula?\[y-y_1=m(x-x_1)\]

OpenStudy (jhannybean):

and your \[m=y'=-2\sqrt{2}\] so all you have to do is use your coordinates and your slope, or \[y'\]

OpenStudy (anonymous):

once you get the derivative, plug in the x value, that gives you the m, then plug in the x value into the original equation and that gives you the y so then you will have x, m, and y, and you can fill in the point slope formula

OpenStudy (anonymous):

Texrbook says: \[y= -2\sqrt{x}+\frac{ \sqrt{2\pi} }{ 2 }-2\sqrt{2}\]

OpenStudy (jhannybean):

That's when you get all your values and plug them in, you will get that as your result.

OpenStudy (anonymous):

im not getting that tho ..

OpenStudy (anonymous):

I dont get where the 1st term is coming from

OpenStudy (e.mccormick):

I see your point on that. I would get \(-2x\sqrt{2}\)

OpenStudy (e.mccormick):

They did something similar with the pi.

OpenStudy (anonymous):

y=-4 sinx \[when x= \frac{ \pi }{ 4 } ,y=-4\sin \frac{ \pi }{ 4 },y=-4\times \frac{ 1 }{ \sqrt{2} }=\frac{ -4\sqrt{2} }{2}=-2\sqrt{2}\] dy/dx=-4cos x \[at x=\frac{ \pi }{ 4 },\frac{ dy }{ dx }=-4\cos \frac{ \pi }{ 4 }=-4*\frac{ 1 }{ \sqrt{2} }=-2\sqrt{2}\] \[eq. of tangent line is y-\left( -2\sqrt{2} \right)=-2\sqrt{2}\left( x-\frac{ \pi }{4} \right)\]

OpenStudy (anonymous):

simplify & get the reqd. eq

OpenStudy (e.mccormick):

oops, yah, went back to the original, not the derivative.

OpenStudy (e.mccormick):

hmmm... but still, how would the x and \(\pi\) end up inside the roots to get the book answer.

OpenStudy (jhannybean):

Shouldn't the slope be \[-2\sqrt{2}\]? His textbook answer is only stating -2, disregarding the rad2. How is that possible?

OpenStudy (jhannybean):

\[y=-2\sqrt{x}+\frac{ \sqrt{2\pi} }{ 2 }-2\sqrt{2}\] is his answer, but i think it should be \[y=-2\sqrt{2}x+\frac{ \sqrt{2\pi} }{2 }-2\sqrt{2}\]

OpenStudy (jhannybean):

and how i got this is by following these steps: \[y=-4sinx\] when \[x=\frac{ \pi }{ 4 }\] So you have to take the derivative to find the slope first. \[y' = -4cosx\] and then you can plug in yourx-value to find y'\[y' = -4\cos(\frac{ \pi }{ 4 }) = -4(\frac{ \sqrt{2} }{ 2 })= -2\sqrt{2}\] And then to find the y-intercept, plug the x-value given into the equation:\[y=-4\sin (\frac{ \pi }{ 4 })= -4(\frac{ \sqrt{2} }{ 2 })=-2\sqrt{2}\] Now you have your slope, the y=-intercept, and your x-intercept. You can use the point slope form to evaluate your equation: \[y-y _{1}=m(x-x _{1})\] INputting values, you get:\[y-(-2\sqrt{2})=-2\sqrt{2}(x-\frac{ \pi }{ 4 })\]\[y+2\sqrt{2}=-2\sqrt{2}x+\frac{ 2\sqrt{2\pi} }{ 4}\]\[y=-2\sqrt{2}x+\frac{ \sqrt{2\pi} }{ 2 }-2\sqrt{2}\]

OpenStudy (jhannybean):

I hope it helps somewhat :\

OpenStudy (anonymous):

pi is not in the square root

OpenStudy (jhannybean):

is it just \[\frac{ \sqrt{2} }{ 2 }\pi\] ?

OpenStudy (e.mccormick):

Exacty, so either he has copied his textbook incorrectly or there is an error in his textbook.